OpenStudy (anonymous):

Give the 5th, and 10th term for the sequence formula below. tn = 1 + 6n 30, 60 31, 61 15, 16 11, 16

OpenStudy (anonymous):

answer only plzz

OpenStudy (anonymous):

just sub 5 and 10 in the eq

OpenStudy (anonymous):

Eg: tn = 1 + 6n t5 = 1+ 6*5 = 31 nw find for 10

OpenStudy (anonymous):

@reonrobe just try..it is very easy..))

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

if u r satisifies..plzz...give medal...by clicking the best answer

OpenStudy (anonymous):

<3